OpenStudy (anonymous):

Simplify completely x^2-100/x^2+2X-120 DIVIDED BY 6x+60/x-2

sam (.sam.):

Factor out the GCF of 6 from each term in the polynomial. ((1)/(6(x)+6(10))*(x^(2)-100)/(x^(2)+2x-120))/(x-2) Factor out the GCF of 6 from 6x+60. ((1)/(6(x+10))*(x^(2)-100)/(x^(2)+2x-120))/(x-2) the binomial can be factored using the difference of squares formula, because both terms are perfect squares. The difference of squares formula is a^(2)-b^(2)=(a-b)(a+b). (((1)/(6(x+10))*((x-10)(x+10))/(x^(2)+2x-120)))/(x-2) In this problem 12*-10=-120 and 12-10=2, so insert 12 as the right hand term of one factor and -10 as the right-hand term of the other factor ((1)/(6(x+10)*((x-10)(x+10))/((x+12)(x-10))))/(x-2) Reduce the expression by canceling out the common factor of (x-10) from the numerator and denominator (((1)/(6(x+10))*(1(x+10))/(x+12)))/(x-2) Cancel the common factor of (x+10) from the denominator of the first expression and the numerator of the second expression ((1)/(6)*(1)/(x+12))/(x-2) Multiply the rational expressions to get (1)/(6(x+12)) ((1)/(6(x+12)))/(x-2) Multiply the rational expressions to get (1)/(6(x-2)(x+12))

OpenStudy (anonymous):

THANK YOU